the sum of a number and it's square is 72. find the numbers

Answer by mananth(16949)

let the number be x
the square of it is x^2
x+x^2=72
x^2+x-72=0
x^2+9x-8x-72=0
x(x+9)-8(x+9)=0
(x+9)(x-8)=0
x= -9 or 8